The volume of a cube is given by V = s3, where s is the length of a side. Find the length of a side of a cube if the Volume is 1000 cm3.
Answer by ptaylor(1894)

V=s^3
V=1000 cm^3 So:
s^3=1000 cm^3 take the cube root of both sides
s=cube root (1000 cm^3)
How do we find the cube root of 1000? We can use a calculator or we can do it by trial and error. Lets do it by trial and error
What number multiplied by itself three times yields 1000?
Is it 5? 5*5*5=25*5=125 ---------------No. Much too small
Is it 9? 9*9*9=81*9=729----------------No. 9 is still too small
Is it 12? 12*12*12=144*12=1728----------No. 12 is too large
So now we know it is somewhere between 9 and 12
How about 10? 10*10*10=100*10=1000--------------BINGO
Then s=10 cm----------------------------------------------ans
Hope this helps---ptaylor